A Bibliography of Publications in Computer Language Magazine

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

26 November 2024 Version 1.07

Title word cross-reference

*PLUS [Har91].

2 [Bis91, Ive90, Raj92, Rib90]. 21st [Tou90].

5.0a [Bin92c].

77/386 [Zei92].

’90s [Keu92b].

Abstract [For92]. Academy [Mur93g]. accuracy [Och92b]. Act [O’B90c]. Action [Nar90]. Actor [Cut92, Ent91]. Ada [Han90]. Add [Kah91]. Add-In [Kah91]. Advanced [Han90, Int90]. Ages [Gil90]. Aid [Haz91]. Ain’t [Cre92]. Algorithm [Rey93]. Algorithms [Sch92a]. almost [Tay86]. Along [Gla90]. also [Och92a].

Among [Sch92c, Hil90g]. Analyst [Keu92d]. ANSI [Rab93, Sak90a]. APLsub [Har91]. Application [Keu91a, Mor90a]. Applications [Raj92]. Applying [Atw90]. Approximating [Pla91a]. Aranda [And90a]. Architect [Keu92c, Keu91b]. Architecting [Raj92]. Architecture [Mor90a]. arithmetic [Och89b, Pla88]. Art [Swa91b, Par90b]. Assembler [Tou90]. Assembling [Gla91]. Assembly [Kyl91b]. Asynchronous [Kiy90b]. Automating [Mud90]. Awk [Ang92, Gil90].

B [Pow90]. B-Tree [Pow90]. baked [Par90c]. Balancing [Och92b]. based [Rie91a]. BASIC [Ang90b]. Basics [Kne92, Win90]. Basing [Wel92]. Behind [Fro90]. Berlin [Zei91a]. Best [Gil90, Hil91a]. Better [Eri93]. Beyond [Eri93, Mis90, Win90]. bigger [Mur92f]. bigots [Hil90g]. Bits [Mor90b, Bak92].

Blistering [Bak93]. Blithering [Bak93]. Blocks [Och92c, Och92d, Och92e, Och93a, Och93b, Och93c, Och93d]. Bones [Gig90]. Booch [Boo92a, Boo92b]. Borland [Buz93, Wag93, Wee90]. Bringing [Tur92]. Brook [Pow92b]. Browser [Ent90a]. Builder [Zei91b]. Building [Och92c, Och92d, Och92e, Och93a, Och93b, Och93c, Och93d, Plu91]. bust [Bak92].

C [Ano88, Bin92c, Car92, Coh92b, Cra93, Dlu90b, Dlu90c, Eck92a, Eck92b, Ent90a, For92, Hel92, Hil90g, Hin92, Hym90, Lan93, LD92, Mud90, Nar90, O’B90e, O’B91a, O’B91i, O’B92i, Par90a, Rei92, Sak90a, Sak90b, TH90, WP91, Wil90]. Candidate [Bul91]. Capturing [Sti90a]. Case [Par91b, Keu92e, PJ93, Par90b, TH90]. CASEing [Keu92a]. Catch [LD92]. Century [Tou90]. Chameleon [Cra93]. Changes [Cal92]. Charming [Far91]. cheap [Mur93h]. Check [Dlu90a]. Check-Out [Dlu90a]. Choreographer [Dav90a]. Class [Cra93, O’B92b, Ber90a, Hop91, Mur92e]. Classy [Wee90]. Clean [Keu92b]. Client [Byr91]. Client/Server [Byr91]. Clipping [Rey92, Rey93]. Close [Keu92c, Keu92d]. Club [Gro88]. COBOL [Bin92a, Can91, DH92, Mis90, Mis91, Rab93, Rie91b]. Code [Ano88, Edg91, Sha92b, Wee92, BT86, Mul91b, Plu93]. Colors [JR90, Pom90]. Command [Ive90]. Command-Line [Ive90]. Common [Ste86]. CommonBase [Vea92]. competition [Gla91]. Compilers [Nar90, Sak90b, Cre90]. comprehension [Sha92b]. Compression [Kyl90a]. Conformance [Plu93]. Connection [Ste86]. Control [Eck91]. Controlling [JR90]. Cooking [Can91]. cooperative [Ano92]. Copyright [Whi93]. Corp [O’B92i]. Cosmetic [And91a, Dav91]. Counting [Eck92b]. Course [Cut91]. cpp [Haz91]. Crafting [Rod90]. Creating [Rey92].

Crunching [Bak93]. Culture [NK92].

Cweb [Lev93].

Data [Duf91, For92, Hel91, KKL^{+}93, Wel92]. Databases [Atw90, Mur92a, Rei93]. Date [BT90]. Debatable [KKL^{+}93]. Debuggers [Kyl92]. Debugging [Dlu90b, Int90, Mud90]. Decisions [Con92e]. Declarations [Mar91]. Design [Cor90, NK92, Ree92, SM90, TH90, WP91, McK91]. Desktop [Tur92]. Determining [Wal93]. Development [Ang90b, Par90a, Kiy90a, Kiy91a, Kiy91b]. Diagrams [Rov90]. Dialogues [Far91]. Difficult [Plu91]. Dimensional [Kra90]. Discovering [O’B92e]. Dither [Lam91]. Documentation [Par91a]. Dog [Gig91]. Doing [Sch92a]. Done [Wee92]. Dusting [Mis91].

Easy [Eck92a, Mar91]. ECOMP [Kyl90a]. Economizing [Pla91b]. edge [Ano92]. Editor [Ano92, Hil90f, Hil90g, Hil90h, Hil90a, Hil90b, Hil90c, Hil90d, Hil90e, Mul91a, O’B90c, O’B90d, O’B90a, O’B90b, O’B91b, O’B91c, O’B91d, O’B91e, O’B91f, O’B91g, O’B91h, O’B92e, O’B92c, O’B92d, O’B93a, O’B93b, O’B93c, O’B93d, O’B93e, O’B93f]. editors [Mul91b]. Eiffel [O’B92g]. Eiffel/S [O’B92g]. Elegance [FG92]. Eleven [Hoo92a]. Encapsulating [Wee90]. End [McT91]. Engineering [O’B92g, Swa91b, Wie93, Keu90, Pla92e]. Engineeringin [Rie91b]. Enhanced [McC90]. Entered [Abe93]. Enterprise [Kar92]. Entities [BB92, Keu92a]. Environment [Suy90]. Environments [Mur93a, Par90a]. Error [Eck92a, LD92]. Events [Kiy90b]. ever [Mul91b]. Evolution [Car92]. Exception [LD92]. Exotic [Gro88]. Expert [Tur92, Lew90]. Exploratorium [Nka92]. Expressions [Pak91]. Extender [Tay90b]. Extensibility [Con90]. EZ [Ano91]. EZ-Install [Ano91]. face [And91c]. face-off [And91c]. faces [Ber90b]. Facts [Hil90f]. Fancy [Rav90]. far [Zei91a]. Fast [And93, McG93, OH91, Sch92b]. fever [Cre90]. Few [Pom90, Spe91]. fictions [Hil90f, Och91a]. Filer [Pow90]. fill [Rey92]. Final [Hil92]. Find [Kyl91a]. Floating [Och89a, Och89b, Pla88]. Floating-point [Och89a, Pla88]. Flying [Gla90]. forces [O’B92k]. Formal [Keu92b, SW92]. Forth [Spe91]. Forths [Spe92]. FORTRAN [Cre90, Zei92]. foundation [Och91c, Och91d]. Four [Cor90, Spe92]. Friends [Sch92c]. function [Rey92]. functional [BT91]. Functions [Con90, Pla91a].

Geller [Ang90a]. Gender [Pak91]. Generalized [Rie91a]. Generation [Cel90]. generators [And91c, Keu91a]. Generics [Han90]. Get [Dav90b, Fro90, OH91, Hoo92a]. Getting [Duf91, Rog90]. global [BT86]. globally [O’B90c]. goes [Rib90]. Good [Pom90, Tyl90]. Goto [Coo91]. Gotta [Cha91]. Graphics [Kra90, Mur92a, Wee90]. Group [Wie93, Zei92]. Groupware [Keu90]. Guessing [Pak91]. GUI [Dav90b, NK92, OH91].

half [Par90c]. half-baked [Par90c]. hammer [Mur92f]. Handling [LD92]. Hanging [Phi90]. Hard [Plu91]. Harvey [Och92f]. Hash [Hel91]. Heap [Nel90]. Hello [Bin92b, Mur92b, Mur92c, Mur92d, Mur93b, Mur93c, Mur93d, Mur93e, Mur93f, O’B92f, O’B93g]. High [Cre90]. High-powered [Cre90]. Hodgman [Rey93]. Hold [Par91a]. Hungarian [Can91].

IBM [And91b]. I’d [Mur93g]. II [Boo92b]. Image [Kyl90a]. Implementing [Wie93]. Inc. [Zei92]. independent [BT86]. Information [Byr91]. Input [Sti90a].

Install [Ano91]. Installer [And91b]. Instant [Bin92c]. Interactive [O’B92g]. Interface [Cor90, Ent90b, Wee90]. intermediate [BT86]. International [Wag93]. Interviews [Est90, KW85]. Introduction [Koo93]. Issues [O’B93h].

Jake [Ang90a]. Jeopardy [Hil92]. Journey [Lee91]. Just [Ang92, Gig91, O’B90e, Plu91].

Ken [KW85]. Keyboard [McC90]. King [Hin92]. know [Mul91b]. Knowledge [BB92].

laid [Hoo92a]. Language

[Ang92, BT86, Gro88, Skj84]. Language- [BT86]. Languages [Rie91a]. Latest [Rab93]. Least [Duf91]. Lee [Och92f]. lex [Lev92]. libraries [Dav90c, Mur92e, O’B92b]. Library [Dlu90a, Ent90b]. Life [Hil90g]. Like [Bro91, Mur93g, Bra93a, Gro88, O’B90d]. limits [Bak92]. Line [Ive90]. Lines [Rav90]. LISP [Ste86]. Literate [Hym90, Lev93]. Logic [BT91]. Logical [O’B92h]. Looking [Mur93h]. Lotus [Kah91].

m1 [Ben90]. Mac [Tay90c]. MacApp [And90b]. Machine [Ste86, BT86]. machine-independent [BT86]. Macintosh [And91c]. Macro [Ben90]. Macros [Mud90]. Made [Kep93, Mar91]. Mainframes [Mis90]. Maintenance [DH92]. make [Hon91]. Making [Hel91, Keu90]. Management [BT90, Fra91]. Manager [Raj92]. Managing [Nel90]. Manifesto [PJ93]. Marriage [Cre91]. match [Tay86]. Matching [Le 91]. Math [McG93, Mur92e]. Matrix [Sch92b]. Me [Fro90, Gil90]. Member [Dlu90c]. Menus [Hil91b]. Metaphone [Phi90]. Method [Boo92a, Boo92b]. Methods [Keu92b, Rie91a]. Metrowerks [Tay90a]. Mice [Hil91b]. Microsoft [Ang90b, Cut91, O’B92i]. Migration [Gla90]. Miles [Lee91]. Minding [Eck92b]. Mini [Ben90, Ent90a]. Mini-Browser [Ent90a]. Modeling [PJCW90]. Modes [Whi90]. Modula [Tay90a, Bis91, Rib90]. Modula-2 [Tay90a, Bis91, Rib90]. Mom [Hil91b]. Monster [DH92]. Month [Gro88]. MS [Sak90b]. MS-DOS [Sak90b]. Multiplies [Sch92b]. Multitasking [Bis91]. Mystery [McT91].

n [Men91]. NASA [Och92f]. need [Mur92f]. no [Par90c]. Notation [Boo92a, Boo92b, PJCW90]. Notes [Ano92, Hil90f, Hil90g, Hil90h, Hil90a, Hil90b, Hil90c, Hil90d, Hil90e, Mul91a, O’B90c, O’B90d, O’B90a, O’B90b, O’B91b, O’B91c, O’B91d, O’B91e, O’B91f, O’B91g, O’B91h, O’B92e, O’B92c, O’B92d, O’B93a, O’B93b, O’B93c, O’B93d, O’B93e, O’B93f]. Numbers [Bak93]. Numeric [Och91a]. Numerics [Och91b].

Object [Atw90, GR90, Hil90h, Nka92, PJCW90, Ree92, Rei93, Rie91a, TH90, WP91, McK91]. Object-based [Rie91a]. Object-Oriented [GR90, PJCW90, TH90, WP91, Ree92, Rei93, McK91]. Objects [Bul91, BB92, Cal92, Con90, Eze90]. ObjectVision [Wag93]. Occam [Skj84]. off [And91c, Hoo92a, Mis91]. Only [O’B92h]. operating [Par90c]. Operator [Rei92]. Opportunities [Rab92]. optimization [BT86]. Optimizations [Lan93]. Optimizing [Nar90]. Oriented [GR90, PJCW90, TH90, WP91, McK91, Ree92, Rei93]. Orthogonality [Ree92]. OS/2 [Ive90, Raj92, Rib90]. Oswald [Och92f]. other [Och91a]. Overloading [Rei92].

Paradigm [Atw90]. parallel [Skj84]. Part [And91a, Boo92a, Boo92b, Dav91, Kiy91b].

Pascal [Pow92b, Vea91b, Buz93, Elk92, Ent90a, Gro88, Kra90, Vea91a]. Pascal-like [Gro88]. Path [Gla90]. PC [Ano88, Par90b]. PCs [And91b]. Peopleware [Con92a, Con92b, Con92c, Con92d, Con93a, Con93b, Con93c, Con93d, Con93e, Con93f]. Perfect [Cre91]. Periscope [Sti90b]. Perl [Sha92a]. Permutation [Cel90]. Philosophy [Kyl90b, KW85]. Phoolishness [Kyl90b]. Pigments [Kiy90a]. PIMs [Fra91]. Planning [Mor90a]. Plug [Men91]. POET [O’B91i]. point [Och89a, Och89b, Pla88]. Pointers [Dlu90c]. Polygon [Rey92, Rey93]. polygon-fill [Rey92]. polynomials [Pla91b]. Pop [Dav90c]. Pop-up [Dav90c]. Popkin [Keu91b]. Portability [Haz91]. Portable [Swa91a, Plu93]. PostScript [Koo93]. Potential [Coh92a]. powered [Cre90]. practice [Och89a, Och89b]. Praise [Hil91a]. Preprocessor [Mud90]. Presentation [Raj92]. Principles [Cor90]. Probabilistic [Sch92a]. processing [Skj84]. Processor [Ben90]. Profession [Hoo92b, Hoo92c, Hoo93a, Hoo93b, Hoo93c]. Professional [Ang90b, Tay90b, Cut92, Vea91a]. Program [Con90]. Programmer [Tay90b, Car92]. Programmers [Sha92a]. Programming [Ber90a, Hoo92b, Hoo92c, Hoo93a, Hoo93b, Hoo93c, KW85, Lev93, O’B93h, Pla92a, Pla92b, Pla92c, Pla92d, Pla93a, Pla93b, Pla93c, Pla93d, Pla93e, Pla93f, Sch92a, Suy90, Dav90c, McK91, Rib90]. Programs [Kyl90b, And91b]. Prograph [Mod90]. Project [Fra91]. Prolcon [Tay90c]. PROLOG [Byr91]. Properties [Pla88]. protect [Men91]. PSE [Tay90a]. Pull [CK91]. Purpose [Pla92a, Pla92b, Pla92c, Pla92d, Pla93a, Pla93b, Pla93c, Pla93d, Pla93e, Pla93f]. Puzzle [Gra90].

Quality [Wal93]. Quibbles [Mor90b].

QuickNet [O’B92j].

Random [Kep93]. Randomly [Sch92a]. re [Keu90]. re-engineering [Keu90]. Ready [Nar90]. Realities [Swa91a]. Realizer [Pow92a]. Realizing [Coh92a]. Recursion [Rex92]. Recursive [SM90]. Reengineering [Rab92]. Reference [Eck92b]. Refining [Bul91]. Regular [Pak91]. Remedial [Pla92e]. Rendering [And93]. Reporting [Eck92a]. representations [Och91a]. rest [Och91b]. Retriever [Ive90]. returns [Bin92a]. Reusable [Eze90]. Reverse [Swa91b]. Revisited [Rex92]. Rexx [Gig91]. right [Och92a]. Robust [Le 91]. Room [Keu92b]. rose [Bra93a]. rotten [Och91c, Och91d]. Rovira [Rov90].

S [O’B92g]. Say [O’B90e]. scaffolds [Kiy90a]. Scheme [Dic92]. Scratchpad [BT91]. Search [FG92]. Searching [Wil90]. Secrets [Sch92c]. self [Tyl90]. sense [Keu90]. sequel [Och91d]. Server [Byr91]. Shake [Hon91]. Sharing [Sch92c]. Shielded [Edg91]. Shining [Hel92]. Shrink [Par91a]. Simplify [Cal92]. Simply [Kep93]. sine [Och92a]. Skin [Gig90]. Small [Wie93]. Smalltalk [Ber90b, Suy90, Zei91b]. Smalltalk/V [Zei91b]. Smart [Wil90]. Smelling [Bra93a]. Software [Fra91, O’B92g, Rie91b, Swa91a, Wal93, Whi93, Wie93, Och92f, Pla92e]. Solving [Gra90]. Son [Och91d]. Sorting [McT91]. Soup [Hil90h]. source [Kyl91a]. Spawn [Fro90]. speed [Och92b]. SpellCode [Ang90a]. Spontaneous [Kyl91b]. SQL [Coh92a]. Standard [Plu91, Rab93, Sak90a]. Started [Rog90]. state [Par90b]. STDs [Rod90]. Stony [Pow92b]. Store [Eck92b]. Strategies [Wel92]. Stretching [Bak92]. String [Le 91]. Structured [Est90, WP91]. STSC [Har91]. stuff [O’B90d]. Style [Cha91, NK92]. Substance [NK92].

Success [Dic92]. sunglasses [Mur93h]. Support [McC90, Plu93]. surgery [And91a, Dav91]. Sutherland [Rey93]. Sutherland-Hodgman [Rey93]. System [Keu91b, Keu92c, Par90c, Ang90b]. Systems [Byr91, Eri93, PJCW90, Tur92].

Table [McG93]. Takara [KW85]. Taking [Eck91]. Taming [DH92, GR90, Kiy90b]. Task [Eck91]. teach [McK91]. Techniques [Int90]. Technology [GR90, O’B90d]. Testing [Bec91, Plu91]. Text [Whi90]. Thank [Mur93g]. Thee [Fro90]. Theory [Och89b, Och89a]. Things [BB92]. Think [Coh92b]. Thousand [Lee91]. Three [Ber90b, Kra90]. Three-Dimensional [Kra90]. til [Cre92]. Together [CK91]. Tomorrow [Wel92]. Toolbox [Sha92a]. Toolkit [Kah91]. Tools [Keu92f, Keu92g, Keu92h, Keu92i, Keu93a, Keu93b, Keu93c, KM93a, KM93b, KM93c, Kiy90a, Kiy91a, Kiy91b, O’B92b, Sha92b]. TopSpeed [Elk92]. Trackword [Gra90]. Trade [Keu92f, Keu92g, Keu92h, Keu92i, Keu93a, Keu93b, Keu93c, KM93a, KM93b, KM93c]. Transition [Abe93]. Tree [Pow90]. Trilogy [Gro88]. Trio [Par91b]. Truth [BT91]. Truth-functional [BT91]. TSR [Dav90c]. Turbo [Kra90, O’B91i, Vea91b]. Types [For92, Och91a].

UK [Skj84]. Understanding

[Eri93, Rab93]. Uniface [Bra93b]. Uniform [PJCW90]. UNIX [Fro90, Keu92e, Par90c]. Unknown [Whi90]. User [Cor90, Eri93, Est90, Sti90a]. Using [Cal92, Haz91, Lev92, Rab93, TH90]. utopia [O’B90d].

v [Ang90b, Ano91, Bin92c, Bra93b, Buz93, Coh92b, Cut92, Dav90a, Elk92, Har91, Mod90, O’B92g, O’B92i, O’B92j, Pow90, Pow92a, Pow92b, Vea91a, Vea92, Zei91b, Zei92, Zei91b]. Variables [Kep93]. Versatility [Gro88]. version [Bro91]. Video [Cut91]. Visible [Keu92d]. Vision [O’B91i]. Vital [O’B92k]. vsDesigner [Lew90].

Wake [Abe93]. want [McK91]. wanted [Mul91b]. Washing [Pla91c]. watchers [Pla91c]. Watcom [Zei92]. Way [JR90]. ways [Hoo92a]. Where [SW92]. Who [Eck92b]. WinApps [Rod90]. Window [Zei91b]. Windows [Kiy90a, Kiy90b, Kiy91a, Kiy91b, Kne92, Rog90, Vea91b, Zei91b]. Winner [O’B92a]. Wonder [Gig91]. Words [Spe91]. Wordz [Tay86]. Workbench [Keu92d]. World [Bin92b, Eri93, Mur92b, Mur92c, Mur92d, Mur93b, Mur93c, Mur93d, Mur93e, Mur93f, O’B92e, O’B92f, O’B93g, Mur93h]. Wrap [Par91a]. Writing [Eze90, Och92f].

Xlib [JR90].

yacc [Lev92]. Yes [Coo91]. Yesterday [Wel92]. You’ve [Abe93].

Zinc [Ent90b]. Zone [Abe93].

References

Abelow:1993:WYE

[Abe93] Dan Abelow. Wake up! you’ve entered the transition zone. Computer Language Magazine, 10(3): 40–??, March 1993. CODEN COMLEF. ISSN 0749-2839.

Andrews:1990:A

[And90a] Brad Andrews. Aranda. Computer Language Magazine, 7(11):101–??, November 1990. CODEN COM-LEF. ISSN 0749-2839.

Andrews:1990:M

[And90b] R. Bradley Andrews. MacApp. Computer Language Magazine, 7 (12):85–??, December 1990. CO-DEN COMLEF. ISSN 0749-2839.

Andrews:1991:CSP

[And91a] R. Bradley Andrews. Cosmetic surgery, part I. Computer Language Magazine, 8(9):75–??, September 1991. CODEN COM-LEF. ISSN 0749-2839.

Andrews:1991:IPI

[And91b] R. Bradley Andrews. Installer programs for IBM PCs. Computer Language Magazine, 8(1):89– ??, January 1991. CODEN COM-LEF. ISSN 0749-2839.

Andrews:1991:MGF

[And91c] R. Bradley Andrews. Macintosh generators face-off. Computer Language Magazine, 8(10):105–??, October 1991. CODEN COMLEF. ISSN 0749-2839.

Anderson:1993:FR

[And93] Peter G. Anderson. Fast rendering. Computer Language Magazine, 10 (2):44–??, February 1993. CODEN COMLEF. ISSN 0749-2839.

Angus:1990:JGS

[Ang90a] Jeff Angus. Jake Geller’s Spell-Code. Computer Language Magazine, 7(6):105–??, June 1990. CO-DEN COMLEF. ISSN 0749-2839.

Angus:1990:MBP

[Ang90b] Jeff Angus. Microsoft BASIC Professional Development System v. 7.0. Computer Language Magazine, 7(3):112–??, March 1990.

CODEN COMLEF. ISSN 0749-2839.

Angell:1992:AAL

[Ang92] Bob Angell. Awk — just another language? Computer Language Magazine, 9(12):55–??, December 1992. CODEN COMLEF. ISSN 0749-2839.

Anonymous:1988:CCP

[Ano88] Anonymous. C code for the PC.

Computer Language Magazine, 5 (2):52, February 1988. CODEN COMLEF. ISSN 0749-2839. advertisement.

Anonymous:1991:EV

[Ano91] Anonymous. EZ-Install v. 2.12.

Computer Language Magazine, 8 (3):8f–??, March 1991. CODEN COMLEF. ISSN 0749-2839.

Anonymous:1992:ENC

[Ano92] Anonymous. Editor’s notes: The cooperative edge. Computer Language Magazine, 9(9):5–??, September 1992. CODEN COM-LEF. ISSN 0749-2839.

Atwood:1990:AOP

[Atw90] Thomas Atwood. Applying the object paradigm to databases. Computer Language Magazine, 7(9): 36–??, September 1990. CODEN COMLEF. ISSN 0749-2839.

Baker:1992:SLB

[Bak92] Steven Baker. Stretching the limits: 32 bits or bust. Computer Language Magazine, 9(6):73–??, June 1992. CODEN COMLEF. ISSN 0749-2839.

Baker:1993:CNB

[Bak93] Steve Baker. Crunching numbers– blistering or blithering. Computer Language Magazine, 10(5):64–??, May 1993. CODEN COMLEF. ISSN 0749-2839.

Bulman:1992:TOE

[BB92] Dave M. Bulman and Erin Kathleen Bulman. Things, objects, entities, and knowledge. Computer Language Magazine, 9(1):44– ??, January 1992. CODEN COM-LEF. ISSN 0749-2839.

Becker:1991:TT

[Bec91] Pete Becker. Testing, testing. Computer Language Magazine, 8 (4):59–??, April 1991. CODEN COMLEF. ISSN 0749-2839.

Bentley:1990:MMM

[Ben90] Jon Bentley. m1: A mini macro processor. Computer Language Magazine, 7(6):47–61, June 1990. CODEN COMLEF. ISSN 0749-2839.

Bergman:1990:PC

[Ber90a] Noel Bergman. Programming with class. Computer Language Magazine, 7(2):107–??, February 1990. CODEN COMLEF. ISSN 0749-2839.

Bergman:1990:TFS

[Ber90b] Noel Bergman. Three faces of Smalltalk. Computer Language Magazine, 7(4):87–??, April 1990. CODEN COMLEF. ISSN 0749-2839.

Binstock:1992:CR

[Bin92a] Andrew Binstock. COBOL returns. Computer Language Magazine, 9(9):97–??, September 1992. CODEN COMLEF. ISSN 0749-2839.

Binstock:1992:HW

[Bin92b] Andrew Binstock. Hello, world!

Computer Language Magazine, 9 (7):19–??, July 1992. CODEN COMLEF. ISSN 0749-2839.

Binstock:1992:ICV

[Bin92c] Andrew Binstock. Instant C v.

5.0a. Computer Language Magazine, 9(5):87–??, May 1992. CO-DEN COMLEF. ISSN 0749-2839.

Bisignani:1991:MM

[Bis91] Michael Bisignani. Multitasking in Modula-2. Computer Language Magazine, 8(2):53–??, February 1991. CODEN COMLEF. ISSN 0749-2839.

Booch:1992:BMNa

[Boo92a] Grady Booch. The Booch method:

Notation, part I. Computer Language Magazine, 9(9):47–??, September 1992. CODEN COM-LEF. ISSN 0749-2839.

Booch:1992:BMNb

[Boo92b] Grady Booch. The Booch method:

Notation, part II. Computer Language Magazine, 9(10):36–??, October 1992. CODEN COMLEF. ISSN 0749-2839.

Bragg:1993:SLR

[Bra93a] Tom Bragg. Smelling like a rose.

Computer Language Magazine, 10 (3):85–??, March 1993. CODEN COMLEF. ISSN 0749-2839.

Bragg:1993:UV

[Bra93b] Tom Bragg. Uniface v. 5.2. Computer Language Magazine, 10(1): 86–??, January 1993. CODEN COMLEF. ISSN 0749-2839.

Brown:1991:LV

[Bro91] Hugh Brown. Like a version. Computer Language Magazine, 8(8):59– ??, August 1991. CODEN COM-LEF. ISSN 0749-2839.

Bal:1986:LMG

[BT86] Henri E. Bal and Andrew S. Tanenbaum. Language- and machineindependent global optimization on intermediate code. Computer Language Magazine, 11(2): 105–121, 1986. CODEN COM-LEF. ISSN 0749-2839.

Bergman-Terrell:1990:DM

[BT90] Eric Bergman-Terrell. Date management. Computer Language Magazine, 7(12):57–??, December 1990. CODEN COMLEF. ISSN 0749-2839.

Boggia-Togna:1991:STL

[BT91] Gianfranco Boggia-Togna. A scratchpad for truth-functional logic. Computer Language Magazine, 8(12):40–??, December 1991. CODEN COMLEF. ISSN 0749-2839.

Bulman:1991:RCO

[Bul91] David Bulman. Refining candidate objects. Computer Language Magazine, 8(1):30–??, January 1991. CODEN COMLEF. ISSN 0749-2839.

Buzbee:1993:BPV

[Buz93] Bruce Buzbee. Borland Pascal v. 7.0. Computer Language Magazine, 10(3):93–??, March 1993. CODEN COMLEF. ISSN 0749-2839.

Byrd:1991:PCS

[Byr91] Lawrence Byrd. PROLOG and client/server information systems. Computer Language Magazine, 8 (3):36–??, March 1991. CODEN COMLEF. ISSN 0749-2839.

Callahan:1992:UOS

[Cal92] Gene Callahan. Using objects to simplify changes. Computer Language Magazine, 9(11):63–??, November 1992. CODEN COM-LEF. ISSN 0749-2839.

Cannistra:1991:CHC

[Can91] William J. Cannistra. Cooking with Hungarian COBOL. Computer Language Magazine, 8(9): 61–??, September 1991. CODEN COMLEF. ISSN 0749-2839.

Carter:1992:ECP

[Car92] Everett Carter. The evolution of a C++ programmer. Computer Language Magazine, 9(8):53–??, August 1992. CODEN COMLEF. ISSN 0749-2839.

Celko:1990:PG

[Cel90] Joe Celko. Permutation generation. Computer Language Magazine, 7(12):65–??, December 1990. CODEN COMLEF. ISSN 0749-

Charney:1991:YGS

[Cha91] Reginald B. Charney. You gotta have style. Computer Language Magazine, 8(10):57–??, October 1991. CODEN COMLEF. ISSN 0749-2839.

Cohen:1991:PT

[CK91] Rich Cohen and Warren Keuffel. Pull together! Computer Language Magazine, 8(8):36–??, August 1991. CODEN COMLEF. ISSN 0749-2839.

Cohen:1992:RPS

[Coh92a] Dave Cohen. Realizing the potential of SQL. Computer Language Magazine, 9(3):53–??, March 1992. CODEN COMLEF. ISSN 0749-2839.

Cohen:1992:TCV

[Coh92b] Dennis R. Cohen. Think C v. 5.0.2. Computer Language Magazine, 9(5):88–??, May 1992. CO-DEN COMLEF. ISSN 0749-2839.

Constantine:1990:OFP

[Con90] Larry L. Constantine. Objects, functions, and program extensibility. Computer Language Magazine, 7(1):34–??, January 1990. CODEN COMLEF. ISSN 0749-2839.

Constantine:1992:Pa

[Con92a] Larry Constantine. Peopleware. Computer Language Magazine, 9 (9):120–??, September 1992. CO-DEN COMLEF. ISSN 0749-2839.

Constantine:1992:Pb

[Con92b] Larry Constantine. Peopleware. Computer Language Magazine, 9 (10):112–??, October 1992. CO-DEN COMLEF. ISSN 0749-2839.

Constantine:1992:Pc

[Con92c] Larry Constantine. Peopleware.

Computer Language Magazine, 9 (11):104–??, November 1992. CO-DEN COMLEF. ISSN 0749-2839.

Constantine:1992:Pd

[Con92d] Larry Constantine. Peopleware.

Computer Language Magazine, 9 (12):112–??, December 1992. CO-DEN COMLEF. ISSN 0749-2839.

Constantine:1992:DD

[Con92e] Larry L. Constantine. Decisions, decisions. Computer Language Magazine, 9(3):112–??, March 1992. CODEN COMLEF. ISSN 0749-2839.

Constantine:1993:Pa

[Con93a] Larry Constantine. Peopleware.

Computer Language Magazine, 10 (1):112–??, January 1993. CODEN COMLEF. ISSN 0749-2839.

Constantine:1993:Pb

[Con93b] Larry Constantine. Peopleware.

Computer Language Magazine, 10 (2):112–??, February 1993. CO-DEN COMLEF. ISSN 0749-2839.

Constantine:1993:Pc

[Con93c] Larry Constantine. Peopleware.

Computer Language Magazine, 10 (3):120–??, March 1993. CODEN COMLEF. ISSN 0749-2839.

Constantine:1993:Pd

[Con93d] Larry Constantine. Peopleware.

Computer Language Magazine, 10 (4):112–??, April 1993. CODEN COMLEF. ISSN 0749-2839.

Constantine:1993:Pe

[Con93e] Larry Constantine. Peopleware. Computer Language Magazine, 10 (5):112–??, May 1993. CODEN COMLEF. ISSN 0749-2839.

Constantine:1993:Pf

[Con93f] Larry Constantine. Peopleware. Computer Language Magazine, 10 (6):104–??, June 1993. CODEN COMLEF. ISSN 0749-2839.

Cooper:1991:GYG

[Coo91] Tim Cooper. Goto? yes, goto! Computer Language Magazine, 8 (5):63–??, May 1991. CODEN COMLEF. ISSN 0749-2839.

Cornish:1990:FPU

[Cor90] Merrill Cornish. Four principles of user interface design. Computer Language Magazine, 7(3):67– ??, March 1990. CODEN COM-LEF. ISSN 0749-2839.

Crawford:1993:CCC

[Cra93] John Crawford. A chameleon class in C++. Computer Language Magazine, 10(5):40–??, May 1993. CO-DEN COMLEF. ISSN 0749-2839.

Crenshaw:1990:FFH

[Cre90] Jack Crenshaw. FORTRAN fever: High-powered compilers. Computer Language Magazine, 7(5): 113–??, May 1990. CODEN COM-LEF. ISSN 0749-2839.

Crenshaw:1991:PM

[Cre91] Jack W. Crenshaw. A perfect marriage. Computer Language Magazine, 8(6):54–??, June 1991. CO-DEN COMLEF. ISSN 0749-2839.

Crenshaw:1992:IAT

[Cre92] Jack Crenshaw. It ain’t over til it’s over. Computer Language Magazine, 9(1):59–??, January 1992. CODEN COMLEF. ISSN 0749-2839.

Cutrona:1991:MVC

[Cut91] Louis Cutrona. Microsoft video course. Computer Language Magazine, 8(4):70–??, April 1991. CO-DEN COMLEF. ISSN 0749-2839.

Cutrona:1992:APV

[Cut92] Lou Cutrona. Actor Professional v. 4.0. Computer Language Magazine, 9(4):75–??, April 1992. CO-DEN COMLEF. ISSN 0749-2839.

Davidson:1990:CV

[Dav90a] Mark Davidson. Choreographer v.

1.1. Computer Language Magazine, 7(11):99–??, November 1990. CODEN COMLEF. ISSN 0749-2839.

Davidson:1990:GGI

[Dav90b] Mark Davidson. Get a GUI on it! Computer Language Magazine, 7(8):81–??, August 1990. CODEN COMLEF. ISSN 0749-2839.

Davidson:1990:TLP

[Dav90c] Mark Davidson. TSR libraries:

Pop-up programming. Computer Language Magazine, 7(5):127–??, May 1990. CODEN COMLEF. ISSN 0749-2839.

Davidson:1991:CSP

[Dav91] Mark Davidson. Cosmetic surgery, part 2. Computer Language Magazine, 8(10):87–??, October 1991.

CODEN COMLEF. ISSN 0749-2839.

Dyson-Hudson:1992:TCM

[DH92] Nigel Dyson-Hudson. Taming the COBOL maintenance monster. Computer Language Magazine, 9 (9):40–??, September 1992. CO-DEN COMLEF. ISSN 0749-2839.

Dickey:1992:SS

[Dic92] Kenneth Dickey. Scheme for success. Computer Language Magazine, 9(6):47–??, June 1992. CO-DEN COMLEF. ISSN 0749-2839.

Dlugosz:1990:LC

[Dlu90a] John Dlugosz. Library check-out. Computer Language Magazine, 7 (5):83–??, May 1990. CODEN COMLEF. ISSN 0749-2839.

Dlugosz:1990:DC

[Dlu90b] John M. Dlugosz. Debugging in C++. Computer Language Magazine, 7(2):32–??, February 1990. CODEN COMLEF. ISSN 0749-2839.

Dlugosz:1990:MPC

[Dlu90c] John M. Dlugosz. Member pointers in C++. Computer Language Magazine, 7(1):79–??, January 1990. CODEN COMLEF. ISSN 0749-2839.

Dufort:1991:GLY

[Duf91] Martin Dufort. Getting the least out of your data. Computer Language Magazine, 8(5):45–??, May 1991. CODEN COMLEF. ISSN 0749-2839.

Eckel:1991:TCT

[Eck91] Bruce Eckel. Taking control to task. Computer Language Magazine, 8(2):38–??, February 1991. CODEN COMLEF. ISSN 0749-2839.

Eckel:1992:EER

[Eck92a] Bruce Eckel. Easy error reporting in C++. Computer Language Magazine, 9(2):77–??, February 1992. CODEN COMLEF. ISSN 0749-2839.

Eckel:1992:WMS

[Eck92b] Bruce Eckel. Who’s minding the store? reference counting in C++. Computer Language Magazine, 9 (5):40–??, May 1992. CODEN COMLEF. ISSN 0749-2839.

Edgar:1991:SC

[Edg91] Bob Edgar. Shielded code. Computer Language Magazine, 8(6):65– ??, June 1991. CODEN COMLEF. ISSN 0749-2839.

Elkins:1992:TPV

[Elk92] T. A. Elkins. TopSpeed Pascal v.

3.02. Computer Language Magazine, 9(4):85–??, April 1992. CO-DEN COMLEF. ISSN 0749-2839.

Entsminger:1990:MCP

[Ent90a] Gary Entsminger. A mini-browser for C++ and Pascal. Computer Language Magazine, 7(8):45– ??, August 1990. CODEN COM-LEF. ISSN 0749-2839.

Entsminger:1990:ZIL

[Ent90b] Gary Entsminger. Zinc Interface Library. Computer Language Magazine, 7(12):73–??, December 1990. CODEN COMLEF. ISSN 0749-2839.

Entsminger:1991:A

[Ent91] Gary Entsminger. Actor 3.0. Computer Language Magazine, 8(3):71– ??, March 1991. CODEN COM-LEF. ISSN 0749-2839.

Erickson:1993:BSB

[Eri93] J. David Erickson. Beyond systems: Better understanding the user’s world. Computer Language Magazine, 10(3):51–??, March 1993. CODEN COMLEF. ISSN 0749-2839.

Estvanik:1990:SUI

[Est90] Steve Estvanik. Structured user interviews. Computer Language Magazine, 7(6):77–??, June 1990. CODEN COMLEF. ISSN 0749-2839.

Ezell:1990:WRO

[Eze90] Ben Ezell. Writing reusable objects. Computer Language Magazine, 7(6):34–??, June 1990. CO-DEN COMLEF. ISSN 0749-2839.

Farah:1991:CD

[Far91] Avraham Farah. Charming dialogues. Computer Language Magazine, 8(7):43–??, July 1991. CO-DEN COMLEF. ISSN 0749-2839.

Fisher:1992:SE

[FG92] Jeff Fisher and Dale Gipson. In search of elegance. Computer Language Magazine, 9(11):36–??, November 1992. CODEN COM-LEF. ISSN 0749-2839.

Ford:1992:ADT

[For92] Dan Ford. Abstract data types in C. Computer Language Magazine, 9(5):59–??, May 1992. CO-DEN COMLEF. ISSN 0749-2839.

Frank:1991:SPM

[Fra91] Maurice Frank. Software project management with PIMs. Computer Language Magazine, 8(8):55– ??, August 1991. CODEN COM-LEF. ISSN 0749-2839.

Frost:1990:GTB

[Fro90] Lyle Frost. Get thee behind me, spawn of UNIX! Computer Language Magazine, 7(7):51–??, July 1990. CODEN COMLEF. ISSN 0749-2839.

Giguere:1990:SB

[Gig90] Eric Giguere. Skin and bones.

Computer Language Magazine, 7 (4):43–??, April 1990. CODEN COMLEF. ISSN 0749-2839.

Giguere:1991:RJW

[Gig91] Eric Giguere. Rexx: Not just a wonder dog. Computer Language Magazine, 8(3):45–??, March 1991. CODEN COMLEF. ISSN 0749-2839.

Gilchrist:1990:AAB

[Gil90] Michael J. Gilchrist. Awk of ages, best for me. Computer Language Magazine, 7(1):85–??, January 1990. CODEN COMLEF. ISSN 0749-2839.

Glass:1990:FAM

[Gla90] Brett Glass. Flying along the migration path. Computer Language Magazine, 7(11):38–??, November 1990. CODEN COMLEF. ISSN 0749-2839.

Glass:1991:AC

[Gla91] Brett Glass. Assembling the competition. Computer Language Magazine, 8(11):77–??, November 1991. CODEN COMLEF. ISSN 0749-2839.

Goldberg:1990:TOT

[GR90] Adele Goldberg and Kenneth S. Rubin. Taming object-oriented technology. Computer Language Magazine, 7(10):34–??, October 1990. CODEN COMLEF. ISSN 0749-2839.

Gray:1990:STP

[Gra90] John Gray. Solving the trackword puzzle. Computer Language Magazine, 7(9):71–??, September 1990. CODEN COMLEF. ISSN 0749-2839.

Grogono:1988:ELM

[Gro88] P. Grogono. Exotic language of the month club: More versatility with Pascal-like Trilogy. Computer Language Magazine, 5(4):83–86, April 1988. CODEN COMLEF. ISSN 0749-2839.

Hanson:1990:AGA

[Han90] David Hanson. Advanced generics in Ada. Computer Language Magazine, 7(6):89–??, June 1990. CO-DEN COMLEF. ISSN 0749-2839.

Hardell:1991:SAP

[Har91] Wes Hardell. STSC APLsub *PLUS v. 1.0. Computer Language Magazine, 8(4):80–??, April 1991.

CODEN COMLEF. ISSN 0749-2839.

Hazard:1991:UCA

[Haz91] William P. Hazard. Using cpp to aid portability. Computer Language Magazine, 8(11):49–??, November 1991. CODEN COM-LEF. ISSN 0749-2839.

Heller:1991:MHY

[Hel91] Steve Heller. Making a hash of your data. Computer Language Magazine, 8(8):47–??, August 1991. CODEN COMLEF. ISSN 0749-2839.

Heller:1992:CSC

[Hel92] Steve Heller. From C to shining C. Computer Language Magazine, 9(6):38–??, June 1992. CODEN COMLEF. ISSN 0749-2839.

Hildebrand:1990:ENa

[Hil90a] J. D. Hildebrand. Editor’s notes.

Computer Language Magazine, 7 (1):5–??, January 1990. CODEN COMLEF. ISSN 0749-2839.

Hildebrand:1990:ENb

[Hil90b] J. D. Hildebrand. Editor’s notes.

Computer Language Magazine, 7 (2):5–??, February 1990. CODEN COMLEF. ISSN 0749-2839.

Hildebrand:1990:ENc

[Hil90c] J. D. Hildebrand. Editor’s notes.

Computer Language Magazine, 7 (3):5–??, March 1990. CODEN COMLEF. ISSN 0749-2839.

Hildebrand:1990:ENd

[Hil90d] J. D. Hildebrand. Editor’s notes.

Computer Language Magazine, 7 (4):5–??, April 1990. CODEN COMLEF. ISSN 0749-2839.

Hildebrand:1990:ENe

[Hil90e] J. D. Hildebrand. Editor’s notes. Computer Language Magazine, 7 (5):5–??, May 1990. CODEN COMLEF. ISSN 0749-2839.

Hildebrand:1990:ENF

[Hil90f] J. D. Hildebrand. Editor’s notes facts and fictions. Computer Language Magazine, 7(9):5–??, September 1990. CODEN COM-LEF. ISSN 0749-2839.

Hildebrand:1990:ENL

[Hil90g] J. D. Hildebrand. Editor’s notes life among the C bigots. Computer Language Magazine, 7(7):5– ??, July 1990. CODEN COMLEF. ISSN 0749-2839.

Hildebrand:1990:ENO

[Hil90h] J. D. Hildebrand. Editor’s notes object soup. Computer Language Magazine, 7(6):5–??, June 1990. CODEN COMLEF. ISSN 0749-2839.

Hildebrand:1991:PB

[Hil91a] J. Hildebrand. In praise of 1990’s best. Computer Language Magazine, 8(4):38–??, April 1991. CO-DEN COMLEF. ISSN 0749-2839.

Hildebrand:1991:MMM

[Hil91b] J. D. Hildebrand. Menus, mice, and mom. Computer Language Magazine, 8(9):42–??, September 1991. CODEN COMLEF. ISSN 0749-2839.

Hildebrand:1992:FJ

[Hil92] J. D. Hildebrand. Final jeopardy.

Computer Language Magazine, 9 (12):73–??, December 1992. CO-DEN COMLEF. ISSN 0749-2839.

Hinners:1992:KC

[Hin92] Bonny V. Hinners. King of C++.

Computer Language Magazine, 9 (5):77–??, May 1992. CODEN COMLEF. ISSN 0749-2839.

Honeycutt:1991:SM

[Hon91] Jerry Honeycutt. Shake and make.

Computer Language Magazine, 8 (2):71–??, February 1991. CODEN COMLEF. ISSN 0749-2839.

Hooten:1992:EWG

[Hoo92a] Karen Hooten. Eleven ways to get laid off. Computer Language Magazine, 9(3):105–??, March 1992. CODEN COMLEF. ISSN 0749-2839.

Hooten:1992:PPa

[Hoo92b] Karen Hooten. Programming by profession. Computer Language Magazine, 9(10):105–??, October 1992. CODEN COMLEF. ISSN 0749-2839.

Hooten:1992:PPb

[Hoo92c] Karen Hooten. Programming by profession. Computer Language Magazine, 9(12):93–??, December 1992. CODEN COMLEF. ISSN 0749-2839.

Hooten:1993:PPa

[Hoo93a] Karen Hooten. Programming by profession. Computer Language Magazine, 10(2):105–??, February 1993. CODEN COMLEF. ISSN 0749-2839.

Hooten:1993:PPb

[Hoo93b] Karen Hooten. Programming by profession. Computer Language Magazine, 10(4):105–??, April 1993. CODEN COMLEF. ISSN 0749-2839.

Hooten:1993:PPc

[Hoo93c] Karen Hooten. Programming by profession. Computer Language Magazine, 10(6):99–??, June 1993. CODEN COMLEF. ISSN 0749-2839.

Hopkins:1991:NC

[Hop91] Jon Hopkins. A new class. Computer Language Magazine, 8(10): 79–??, October 1991. CODEN COMLEF. ISSN 0749-2839.

Hyman:1990:LC

[Hym90] Marco S. Hyman. Literate C++. Computer Language Magazine, 7 (7):67–68, 70, 72, 74–77, 79, July 1990. CODEN COMLEF. ISSN 0749-2839.

Intersimone:1990:ADT

[Int90] David Intersimone. Advanced debugging techniques. Computer Language Magazine, 7(2):59–??, February 1990. CODEN COM-LEF. ISSN 0749-2839.

Ives:1990:CRO

[Ive90] William Eric Ives. Commandline retriever for OS/2. Computer Language Magazine, 7(11):65–??, November 1990. CODEN COM-LEF. ISSN 0749-2839.

Johnson:1990:CCX

[JR90] Eric F. Johnson and Kevin Reichard. Controlling colors the Xlib way. Computer Language Magazine, 7(7):34–??, July 1990. CO-DEN COMLEF. ISSN 0749-2839.

Kahn:1991:LAT

[Kah91] Jeffrey Kahn. Lotus Add-In Toolkit 3.1. Computer Language Magazine, 8(4):66–??, April 1991. CODEN COMLEF. ISSN 0749-2839.

Karbowski:1992:IE

[Kar92] Joe Karbowski. E is for enterprise.

Computer Language Magazine, 9 (3):42–??, March 1992. CODEN COMLEF. ISSN 0749-2839.

Keppler:1993:RVM

[Kep93] Karl Keppler. Random variables made simply. Computer Language Magazine, 10(6):67–??, June 1993. CODEN COMLEF. ISSN 0749-2839.

Keuffel:1990:MSI

[Keu90] Warren Keuffel. Making sense of it all: Groupware for re-engineering. Computer Language Magazine, 7 (6):93–??, June 1990. CODEN COMLEF. ISSN 0749-2839.

Keuffel:1991:AG

[Keu91a] Warren Keuffel. Application generators. Computer Language Magazine, 8(9):85–??, September 1991. CODEN COMLEF. ISSN 0749-2839.

Keuffel:1991:PSA

[Keu91b] Warren Keuffel. Popkin System Architect. Computer Language Magazine, 8(1):92–??, January 1991. CODEN COMLEF. ISSN 0749-2839.

Keuffel:1992:CE

[Keu92a] Warren Keuffel. CASEing the entities. Computer Language Magazine, 9(3):29–??, March 1992. CO-DEN COMLEF. ISSN 0749-2839.

Keuffel:1992:CYR

[Keu92b] Warren Keuffel. Clean your room: Formal methods in the ’90s. Computer Language Magazine, 9(7):38– ??, July 1992. CODEN COMLEF. ISSN 0749-2839.

Keuffel:1992:CSA

[Keu92c] Warren Keuffel. Close up: System architect. Computer Language Magazine, 9(4):29–??, April 1992. CODEN COMLEF. ISSN 0749-2839.

Keuffel:1992:CVA

[Keu92d] Warren Keuffel. Close up: Visible analyst workbench. Computer Language Magazine, 9(5):35– ??, May 1992. CODEN COMLEF. ISSN 0749-2839.

Keuffel:1992:CU

[Keu92e] Warren Keuffel. In CASE of UNIX. Computer Language Magazine, 9 (1):69–??, January 1992. CODEN COMLEF. ISSN 0749-2839.

Keuffel:1992:TTa

[Keu92f] Warren Keuffel. Tools of the trade. Computer Language Magazine, 9 (9):31–??, September 1992. CO-DEN COMLEF. ISSN 0749-2839.

Keuffel:1992:TTb

[Keu92g] Warren Keuffel. Tools of the trade.

Computer Language Magazine, 9 (10):31–??, October 1992. CODEN COMLEF. ISSN 0749-2839.

Keuffel:1992:TTc

[Keu92h] Warren Keuffel. Tools of the trade.

Computer Language Magazine, 9 (11):31–??, November 1992. CO-DEN COMLEF. ISSN 0749-2839.

Keuffel:1992:TTd

[Keu92i] Warren Keuffel. Tools of the trade.

Computer Language Magazine, 9 (12):33–??, December 1992. CO-DEN COMLEF. ISSN 0749-2839.

Keuffel:1993:TTa

[Keu93a] Warren Keuffel. Tools of the trade.

Computer Language Magazine, 10 (1):35–??, January 1993. CODEN COMLEF. ISSN 0749-2839.

Keuffel:1993:TTb

[Keu93b] Warren Keuffel. Tools of the trade.

Computer Language Magazine, 10 (2):35–??, February 1993. CODEN COMLEF. ISSN 0749-2839.

Keuffel:1993:TTc

[Keu93c] Warren Keuffel. Tools of the trade.

Computer Language Magazine, 10 (3):31–??, March 1993. CODEN COMLEF. ISSN 0749-2839.

Kiyooka:1990:PSW

[Kiy90a] Gen Kiyooka. Pigments and scaffolds: Windows development tools. Computer Language Magazine, 7 (3):91–??, March 1990. CODEN COMLEF. ISSN 0749-2839.

Kiyooka:1990:TAE

[Kiy90b] Gen Kiyooka. Taming asynchronous events under Windows. Computer Language Magazine, 7 (3):34–??, March 1990. CODEN COMLEF. ISSN 0749-2839.

Kiyooka:1991:WDTa

[Kiy91a] Gen Kiyooka. Windows development tools. Computer Language Magazine, 8(2):87–??, February 1991. CODEN COMLEF. ISSN 0749-2839.

Kiyooka:1991:WDTb

[Kiy91b] Gen Kiyooka. Windows development tools: Part 2. Computer Language Magazine, 8(3):59– ??, March 1991. CODEN COM-LEF. ISSN 0749-2839.

Katz:1993:DD

[KKL^{+}93] Peter Katz, John Kornatowski, Mary Loomis, Anu Shukla, and Larry O’Brien. Debatable data. Computer Language Magazine, 10 (1):55–??, January 1993. CODEN COMLEF. ISSN 0749-2839.

Keuffel:1993:TTd

[KM93a] Warren Keuffel and John Merk. Tools of the trade. Computer Language Magazine, 10(4):31–??, April 1993. CODEN COMLEF. ISSN 0749-2839.

Keuffel:1993:TTe

[KM93b] Warren Keuffel and John Merk. Tools of the trade. Computer Language Magazine, 10(5):31–??, May 1993. CODEN COMLEF. ISSN 0749-2839.

Keuffel:1993:TTf

[KM93c] Warren Keuffel and John Merk.

Tools of the trade. Computer Language Magazine, 10(6):31–??, June 1993. CODEN COMLEF. ISSN 0749-2839.

Knecht:1992:BW

[Kne92] Ken Knecht. Basics for Windows.

Computer Language Magazine, 9 (11):71–??, November 1992. CO-DEN COMLEF. ISSN 0749-2839.

Koomey:1993:IP

[Koo93] Gregor Koomey. An introduction to PostScript. Computer Language Magazine, 10(1):67–??, January 1993. CODEN COMLEF. ISSN 0749-2839.

Kraidin:1990:TGT

[Kra90] Jonathan Kraidin. Threedimensional graphics in Turbo Pascal. Computer Language Magazine, 7(9):81–??, September 1990. CO-DEN COMLEF. ISSN 0749-2839.

Knuth:1985:PPI

[KW85] Donald E. Knuth and Niklaus Wirth. Programming philosophy (interviews by Ken Takara). Computer Language Magazine, 2(5):25– 35, May 1985. CODEN COMLEF. ISSN 0749-2839.

Kyle:1990:EIC

[Kyl90a] Jim Kyle. ECOMP image compression. Computer Language Magazine, 7(12):78–??, December 1990. CODEN COMLEF. ISSN 0749-

Kyle:1990:PPP

[Kyl90b] Jim Kyle. Of programs, philosophy, and phoolishness. Computer Language Magazine, 7(1):71– ??, January 1990. CODEN COM-LEF. ISSN 0749-2839.

Kyle:1991:FS

[Kyl91a] Jim Kyle. Find the source. Computer Language Magazine, 8(11): 91–??, November 1991. CODEN COMLEF. ISSN 0749-2839.

Kyle:1991:SA

[Kyl91b] Jim Kyle. Spontaneous assembly. Computer Language Magazine, 8 (3):75–??, March 1991. CODEN COMLEF. ISSN 0749-2839.

Kyle:1992:D

[Kyl92] Jim Kyle. Debuggers. Computer Language Magazine, 9(2):91– ??, February 1992. CODEN COM-LEF. ISSN 0749-2839.

Lamothe:1991:DI

[Lam91] Andre Lamothe. Dither it! Computer Language Magazine, 8(9): 51–??, September 1991. CODEN COMLEF. ISSN 0749-2839.

Lane:1993:OC

[Lan93] Alex Lane. Optimizations in C. Computer Language Magazine, 10 (5):53–??, May 1993. CODEN COMLEF. ISSN 0749-2839.

Leary:1992:CEC

[LD92] Sean Leary and Desmond D’Souza. Catch the error: C++ exception handling. Computer Language Magazine, 9(10):63–??, October 1992. CODEN COMLEF. ISSN 0749-2839.

LeBret:1991:RSM

[Le 91] Christophe Le Bret. Robust string matching. Computer Language Magazine, 8(12):71–??, December 1991. CODEN COMLEF. ISSN 0749-2839.

Lee:1991:JTM

[Lee91] Elgin Lee. The journey of a thousand miles. Computer Language Magazine, 8(10):44–??, October 1991. CODEN COMLEF. ISSN 0749-2839.

Levine:1992:ULY

[Lev92] John R. Levine. Using lex and yacc. Computer Language Magazine, 9(12):67–??, December 1992. CODEN COMLEF. ISSN 0749-2839.

Levy:1993:LPC

[Lev93] Silvio Levy. Literate programming and Cweb. Computer Language Magazine, 10(1):67–68, 70, January 1993. CODEN COMLEF. ISSN 0749-2839.

Lewis:1990:VE

[Lew90] Scott Lewis. vsDesigner Expert.

Computer Language Magazine, 7 (9):95–??, September 1990. CO-DEN COMLEF. ISSN 0749-2839.

Martin:1991:DME

[Mar91] Steven R. Martin. Declarations made easy. Computer Language Magazine, 8(5):71–??, May 1991. CODEN COMLEF. ISSN 0749-2839.

McCoy:1990:EKS

[McC90] Jack E. McCoy. Enhanced keyboard support. Computer Language Magazine, 7(4):77–??, April 1990. CODEN COMLEF. ISSN 0749-2839.

McGrath:1993:FMT

[McG93] Gary McGrath. Fast math and the table. Computer Language Magazine, 10(4):67–??, April 1993. CO-DEN COMLEF. ISSN 0749-2839.

McKim:1991:WWT

[McK91] James C. McKim. Why I want to teach object-oriented programming and design. Computer Language Magazine, 8(9):133–??, September 1991. CODEN COM-LEF. ISSN 0749-2839.

McTiernan:1991:EMS

[McT91] Kelly McTiernan. An end to the mystery of sorting. Computer Language Magazine, 8(7):57–??, July 1991. CODEN COMLEF. ISSN 0749-2839.

Mendelsohn:1991:PP

[Men91] Alex Mendelsohn. Plug ‘n’ protect. Computer Language Magazine, 8(7):75–??, July 1991. CO-DEN COMLEF. ISSN 0749-2839.

Mischel:1990:CBM

[Mis90] Jim Mischel. COBOL: Beyond mainframes. Computer Language Magazine, 7(7):95–??, July 1990. CODEN COMLEF. ISSN 0749-2839.

Mischel:1991:DC

[Mis91] Jim Mischel. Dusting off COBOL. Computer Language Magazine, 8 (11):40–??, November 1991. CO-DEN COMLEF. ISSN 0749-2839.

Modell:1990:PV

[Mod90] Howard Modell. Prograph v. 1.2.

Computer Language Magazine, 7 (2):119–??, February 1990. CO-DEN COMLEF. ISSN 0749-2839.

Moriarty:1990:PAA

[Mor90a] Terry Moriarty. Planning an application architecture. Computer Language Magazine, 7(9):59–??, September 1990. CODEN COM-LEF. ISSN 0749-2839.

Morton:1990:QB

[Mor90b] Mike Morton. Quibbles and bits.

Computer Language Magazine, 7 (12):45–??, December 1990. CO-DEN COMLEF. ISSN 0749-2839.

Mudd:1990:ACD

[Mud90] John Mudd. Automating C debugging with preprocessor macros. Computer Language Magazine, 7 (2):69–??, February 1990. CODEN COMLEF. ISSN 0749-2839.

Mullholland:1991:EN

[Mul91a] Michael Mullholland. Editor’s notes. Computer Language Magazine, 8(5):5–??, May 1991. CO-DEN COMLEF. ISSN 0749-2839.

Mullholland:1991:MTY

[Mul91b] Michael Mullholland. More than you ever wanted to know about code editors. Computer Language Magazine, 8(6):75–??, June 1991. CODEN COMLEF. ISSN 0749-2839.

Murphy:1992:GD

[Mur92a] Thomas Murphy. Graphics databases. Computer Language Magazine, 9(8):67–??, August 1992. CODEN COMLEF. ISSN 0749-2839.

Murphy:1992:HWa

[Mur92b] Thomas Murphy. Hello, world! Computer Language Magazine, 9 (10):17–??, October 1992. CODEN COMLEF. ISSN 0749-2839.

Murphy:1992:HWb

[Mur92c] Thomas Murphy. Hello, world! Computer Language Magazine, 9 (11):17–??, November 1992. CO-DEN COMLEF. ISSN 0749-2839.

Murphy:1992:HWc

[Mur92d] Thomas Murphy. Hello, world! Computer Language Magazine, 9 (12):17–??, December 1992. CO-DEN COMLEF. ISSN 0749-2839.

Murphy:1992:MCL

[Mur92e] Thomas Murphy. Math class libraries. Computer Language Magazine, 9(10):79–??, October 1992. CODEN COMLEF. ISSN 0749-2839.

Murphy:1992:WYN

[Mur92f] Thomas Murphy. When you need a bigger hammer. Computer Language Magazine, 9(7):67–??, July 1992. CODEN COMLEF. ISSN 0749-2839.

Murphy:1993:E

[Mur93a] Thomas Murphy. Environments. Computer Language Magazine, 10 (5):79–??, May 1993. CODEN COMLEF. ISSN 0749-2839.

Murphy:1993:HWa

[Mur93b] Thomas Murphy. Hello, world! Computer Language Magazine, 10 (1):17–??, January 1993. CODEN COMLEF. ISSN 0749-2839.

Murphy:1993:HWb

[Mur93c] Thomas Murphy. Hello, world!

Computer Language Magazine, 10 (2):17–??, February 1993. CODEN COMLEF. ISSN 0749-2839.

Murphy:1993:HWc

[Mur93d] Thomas Murphy. Hello, world!

Computer Language Magazine, 10 (3):17–??, March 1993. CODEN COMLEF. ISSN 0749-2839.

Murphy:1993:HWd

[Mur93e] Thomas Murphy. Hello, world!

Computer Language Magazine, 10 (4):17–??, April 1993. CODEN COMLEF. ISSN 0749-2839.

Murphy:1993:HWe

[Mur93f] Thomas Murphy. Hello, world!

Computer Language Magazine, 10 (6):19–??, June 1993. CODEN COMLEF. ISSN 0749-2839.

Murphy:1993:ILT

[Mur93g] Thomas Murphy. I’d like to thank the academy. Computer Language Magazine, 10(4):36–??, April 1993. CODEN COMLEF. ISSN 0749-2839.

Murphy:1993:LAW

[Mur93h] Thomas Murphy. Looking at the world through cheap sunglasses. Computer Language Magazine, 10 (2):63–??, February 1993. CODEN COMLEF. ISSN 0749-2839.

Naro:1990:OCC

[Nar90] Rick Naro. Optimizing C compilers: Ready for action. Computer Language Magazine, 7(5):47– ??, May 1990. CODEN COMLEF. ISSN 0749-2839.

Nelson:1990:MH

[Nel90] Mark Nelson. Managing the heap. Computer Language Magazine, 7 (2):43–??, February 1990. CODEN COMLEF. ISSN 0749-2839.

Nixdorf:1992:SSG

[NK92] Troy Nixdorf and Gen Kiyooka. Substance and style: GUI design and culture. Computer Language Magazine, 9(2):42–??, February 1992. CODEN COMLEF. ISSN 0749-2839.

Nkanta:1992:OE

[Nka92] Ed Nkanta. Object exploratorium. Computer Language Magazine, 9 (11):55–??, November 1992. CO-DEN COMLEF. ISSN 0749-2839.

OBrien:1990:ENb

[O’B90a] Larry O’Brien. Editor’s notes. Computer Language Magazine, 7 (11):5–??, November 1990. CO-DEN COMLEF. ISSN 0749-2839.

OBrien:1990:ENc

[O’B90b] Larry O’Brien. Editor’s notes. Computer Language Magazine, 7 (12):5–??, December 1990. CO-DEN COMLEF. ISSN 0749-2839.

OBrien:1990:ENA

[O’B90c] Larry O’Brien. Editor’s notes act globally. Computer Language Magazine, 7(8):5–??, August 1990. CO-DEN COMLEF. ISSN 0749-2839.

OBrien:1990:ENT

[O’B90d] Larry O’Brien. Editor’s notes technology, utopia, stuff like that.

Computer Language Magazine, 7 (10):5–??, October 1990. CODEN COMLEF. ISSN 0749-2839.

OBrien:1990:JSC

[O’B90e] Larry O’Brien. Just say C. Computer Language Magazine, 7(5):38– ??, May 1990. CODEN COMLEF. ISSN 0749-2839.

OBrien:1991:CCC

[O’B91a] Larry O’Brien. C, C, C, C. Computer Language Magazine, 8(5):81– ??, May 1991. CODEN COMLEF. ISSN 0749-2839.

OBrien:1991:ENa

[O’B91b] Larry O’Brien. Editor’s notes.

Computer Language Magazine, 8 (1):5–??, January 1991. CODEN COMLEF. ISSN 0749-2839.

OBrien:1991:ENb

[O’B91c] Larry O’Brien. Editor’s notes.

Computer Language Magazine, 8 (2):5–??, February 1991. CODEN COMLEF. ISSN 0749-2839.

OBrien:1991:ENc

[O’B91d] Larry O’Brien. Editor’s notes.

Computer Language Magazine, 8 (3):5–??, March 1991. CODEN COMLEF. ISSN 0749-2839.

OBrien:1991:ENd

[O’B91e] Larry O’Brien. Editor’s notes.

Computer Language Magazine, 8 (4):5–??, April 1991. CODEN COMLEF. ISSN 0749-2839.

OBrien:1991:ENe

[O’B91f] Larry O’Brien. Editor’s notes.

Computer Language Magazine, 8 (6):5–??, June 1991. CODEN COMLEF. ISSN 0749-2839.

OBrien:1991:ENf

[O’B91g] Larry O’Brien. Editor’s notes. Computer Language Magazine, 8 (7):5–??, July 1991. CODEN COMLEF. ISSN 0749-2839.

OBrien:1991:ENg

[O’B91h] Larry O’Brien. Editor’s notes. Computer Language Magazine, 8 (8):5–??, August 1991. CODEN COMLEF. ISSN 0749-2839.

OBrien:1991:TVC

[O’B91i] Larry O’Brien. Turbo Vision for C++ and POET. Computer Language Magazine, 8(9): 19–??, September 1991. CODEN COMLEF. ISSN 0749-2839.

OBrien:1992:WI

[O’B92a] Larry O’Brien. And the winner is. Computer Language Magazine, 9(4):36–??, April 1992. CODEN COMLEF. ISSN 0749-2839.

OBrien:1992:CLT

[O’B92b] Larry O’Brien. Class libraries (tools). Computer Language Magazine, 9(10):83–??, October 1992. CODEN COMLEF. ISSN 0749-2839.

OBrien:1992:ENa

[O’B92c] Larry O’Brien. Editor’s notes. Computer Language Magazine, 9 (11):5–??, November 1992. CO-DEN COMLEF. ISSN 0749-2839.

OBrien:1992:ENb

[O’B92d] Larry O’Brien. Editor’s notes. Computer Language Magazine, 9 (12):5–??, December 1992. CO-DEN COMLEF. ISSN 0749-2839.

OBrien:1992:END

[O’B92e] Larry O’Brien. Editor’s notes: Discovering the new world. Computer Language Magazine, 9(10):5– ??, October 1992. CODEN COM-LEF. ISSN 0749-2839.

OBrien:1992:HW

[O’B92f] Larry O’Brien. Hello, world! Computer Language Magazine, 9(9): 17–??, September 1992. CODEN COMLEF. ISSN 0749-2839.

OBrien:1992:ISE

[O’B92g] Larry O’Brien. Interactive Software Engineering Inc.’s Eiffel/S v. 1.0. Computer Language Magazine, 9(4):19–??, April 1992. CO-DEN COMLEF. ISSN 0749-2839.

OBrien:1992:IOL

[O’B92h] Larry O’Brien. It’s only logical.

Computer Language Magazine, 9 (12):61–??, December 1992. CO-DEN COMLEF. ISSN 0749-2839.

OBrien:1992:MCC

[O’B92i] Larry O’Brien. Microsoft Corp.’s C and C++ v. 7.0. Computer Language Magazine, 9(5):19–??, May 1992. CODEN COMLEF. ISSN 0749-2839.

OBrien:1992:QV

[O’B92j] Larry O’Brien. QuickNet v. 1.0.

Computer Language Magazine, 9 (3):79–??, March 1992. CODEN COMLEF. ISSN 0749-2839.

OBrien:1992:VF

[O’B92k] Larry O’Brien. Vital forces. Computer Language Magazine, 9(3):5– ??, March 1992. CODEN COM-LEF. ISSN 0749-2839.

OBrien:1993:ENa

[O’B93a] Larry O’Brien. Editor’s notes. Computer Language Magazine, 10 (1):5–??, January 1993. CODEN COMLEF. ISSN 0749-2839.

OBrien:1993:ENb

[O’B93b] Larry O’Brien. Editor’s notes. Computer Language Magazine, 10 (2):5–??, February 1993. CODEN COMLEF. ISSN 0749-2839.

OBrien:1993:ENc

[O’B93c] Larry O’Brien. Editor’s notes. Computer Language Magazine, 10 (3):5–??, March 1993. CODEN COMLEF. ISSN 0749-2839.

OBrien:1993:ENd

[O’B93d] Larry O’Brien. Editor’s notes. Computer Language Magazine, 10 (4):5–??, April 1993. CODEN COMLEF. ISSN 0749-2839.

OBrien:1993:ENe

[O’B93e] Larry O’Brien. Editor’s notes. Computer Language Magazine, 10 (5):7–??, May 1993. CODEN COMLEF. ISSN 0749-2839.

OBrien:1993:ENf

[O’B93f] Larry O’Brien. Editor’s notes. Computer Language Magazine, 10 (6):7–??, June 1993. CODEN COMLEF. ISSN 0749-2839.

OBrien:1993:HW

[O’B93g] Larry O’Brien. Hello, world! Computer Language Magazine, 10(5): 19–??, May 1993. CODEN COM-LEF. ISSN 0749-2839.

OBrien:1993:IP

[O’B93h] Larry O’Brien. Issues of programming. Computer Language Magazine, 10(1):44–??, January 1993. CODEN COMLEF. ISSN 0749-2839.

Ochs:1989:FPT

[Och89a] T. Ochs. Floating-point theory and practice. Computer Language Magazine, 6(3):67–81, March 1989. CODEN COMLEF. ISSN 0749-2839.

Ochs:1989:TPF

[Och89b] T. Ochs. Theory and practice (floating point arithmetic). Computer Language Magazine, 6(3):67– 70, 72, 74, 77–78, 80–81, March 1989. CODEN COMLEF. ISSN 0749-2839.

Ochs:1991:NTR

[Och91a] T. Ochs. Numeric types, representations, and other fictions. Computer Language Magazine, 8(8):93– 101, August 1991. CODEN COM-LEF. ISSN 0749-2839.

Ochs:1991:NRU

[Och91b] T. Ochs. Numerics for the rest of us. Computer Language Magazine, 8(10):113–127, October 1991. CO-DEN COMLEF. ISSN 0749-2839.

Ochs:1991:RF

[Och91c] T. Ochs. A rotten foundation.

Computer Language Magazine, 8 (2):103–107, February 1991. CO-DEN COMLEF. ISSN 0749-2839.

Ochs:1991:SRF

[Och91d] T. Ochs. Son of rotten foundation:

The sequel. Computer Language Magazine, 8(3):85–91, March 1991. CODEN COMLEF. ISSN 0749-2839.

Ochs:1992:SIR

[Och92a] T. Ochs. This sine is also right! Computer Language Magazine, 9 (1):89–93, January 1992. CODEN COMLEF. ISSN 0749-2839.

Ochs:1992:BSA

[Och92b] Tom Ochs. Balancing speed and accuracy. Computer Language Magazine, 9(3):87–??, March 1992. CODEN COMLEF. ISSN 0749-2839.

Ochs:1992:BBa

[Och92c] Tom Ochs. Building blocks. Computer Language Magazine, 9(10): 89–??, October 1992. CODEN COMLEF. ISSN 0749-2839.

Ochs:1992:BBb

[Och92d] Tom Ochs. Building blocks. Computer Language Magazine, 9(11): 83–??, November 1992. CODEN COMLEF. ISSN 0749-2839.

Ochs:1992:BBc

[Och92e] Tom Ochs. Building blocks. Computer Language Magazine, 9(12): 83–??, December 1992. CODEN COMLEF. ISSN 0749-2839.

Ochs:1992:ILH

[Och92f] Tom Ochs. Is Lee Harvey Oswald writing NASA’s software? Computer Language Magazine, 9(5):93– ??, May 1992. CODEN COMLEF. ISSN 0749-2839.

Ochs:1993:BBa

[Och93a] Tom Ochs. Building blocks. Computer Language Magazine, 10(1):

105–??, January 1993. CODEN COMLEF. ISSN 0749-2839.

Ochs:1993:BBb

[Och93b] Tom Ochs. Building blocks. Computer Language Magazine, 10(3): 99–??, March 1993. CODEN COMLEF. ISSN 0749-2839.

Ochs:1993:BBc

[Och93c] Tom Ochs. Building blocks. Computer Language Magazine, 10(4): 89–??, April 1993. CODEN COM-LEF. ISSN 0749-2839.

Ochs:1993:BBd

[Och93d] Tom Ochs. Building blocks. Computer Language Magazine, 10(6): 83–??, June 1993. CODEN COM-LEF. ISSN 0749-2839.

Orfali:1991:GGF

[OH91] Robert Orfali and Dan Harkey. Get GUI fast! Computer Language Magazine, 8(7):36–??, July 1991. CODEN COMLEF. ISSN 0749-2839.

Pakin:1991:REG

[Pak91] Scott Pakin. Regular expressions and gender guessing. Computer Language Magazine, 8(12):59–??, December 1991. CODEN COM-LEF. ISSN 0749-2839.

Parker:1990:CDE

[Par90a] Tim Parker. C development environments. Computer Language Magazine, 7(5):97–??, May 1990. CODEN COMLEF. ISSN 0749-2839.

Parker:1990:CPS

[Par90b] Tim Parker. CASE on the PC: The state of the art. Computer Language Magazine, 7(1):89–??, January 1990. CODEN COMLEF. ISSN 0749-2839.

Parker:1990:UIN

[Par90c] Tim Parker. UNIX is no half-baked operating system. Computer Language Magazine, 7(7):83–??, July 1990. CODEN COMLEF. ISSN 0749-2839.

Pariseau:1991:DSW

[Par91a] David Pariseau. Documentation — hold the shrink wrap. Computer Language Magazine, 8(11): 69–??, November 1991. CODEN COMLEF. ISSN 0749-2839.

Parker:1991:CT

[Par91b] Tim Parker. A case trio. Computer Language Magazine, 8(1):73– ??, January 1991. CODEN COM-LEF. ISSN 0749-2839.

Philips:1990:HM

[Phi90] Lawrence Philips. Hanging on the metaphone. Computer Language Magazine, 7(12):38–??, December 1990. CODEN COMLEF. ISSN 0749-2839.

Page-Jones:1993:CM

[PJ93] Meilir Page-Jones. The CASE manifesto. Computer Language Magazine, 10(6):38–??, June 1993. CODEN COMLEF. ISSN 0749-2839.

Page-Jones:1990:MOS

[PJCW90] Meilir Page-Jones, Larry L. Constantine, and Steven Weiss. Modeling object-oriented systems: The uniform object notation. Computer Language Magazine, 7(10):69–??, October 1990. CODEN COMLEF. ISSN 0749-2839.

Plauger:1988:PFP

[Pla88] P. J. Plauger. Properties of floating-point arithmetic. Computer Language Magazine, 5(3):17– 22, March 1988. CODEN COM-LEF. ISSN 0749-2839.

Plauger:1991:AF

[Pla91a] P. J. Plauger. Approximating functions. Computer Language Magazine, 8(6):17–25, June 1991. CO-DEN COMLEF. ISSN 0749-2839.

Plauger:1991:EP

[Pla91b] P. J. Plauger. Economizing polynomials. Computer Language Magazine, 8(7):21–27, July 1991. CO-DEN COMLEF. ISSN 0749-2839.

Plauger:1991:WW

[Pla91c] P. J. Plauger. Washing the watchers. Computer Language Magazine, 8(9):23–32, September 1991. CO-DEN COMLEF. ISSN 0749-2839.

Plauger:1992:PPa

[Pla92a] P. J. Plauger. Programming on purpose. Computer Language Magazine, 9(9):25–??, September 1992. CODEN COMLEF. ISSN 0749-2839.

Plauger:1992:PPb

[Pla92b] P. J. Plauger. Programming on purpose. Computer Language Magazine, 9(10):25–??, October 1992. CODEN COMLEF. ISSN 0749-2839.

Plauger:1992:PPc

[Pla92c] P. J. Plauger. Programming on purpose. Computer Language Magazine, 9(11):25–??, November 1992. CODEN COMLEF. ISSN 0749-2839.

Plauger:1992:PPd

[Pla92d] P. J. Plauger. Programming on purpose. Computer Language Magazine, 9(12):25–??, December 1992. CODEN COMLEF. ISSN 0749-2839.

Plauger:1992:RSE

[Pla92e] P. J. Plauger. Remedial software engineering. Computer Language Magazine, 9(3):23–??, March 1992. CODEN COMLEF. ISSN 0749-2839.

Plauger:1993:PPa

[Pla93a] P. J. Plauger. Programming on purpose. Computer Language Magazine, 10(1):25–??, January 1993. CODEN COMLEF. ISSN 0749-2839.

Plauger:1993:PPb

[Pla93b] P. J. Plauger. Programming on purpose. Computer Language Magazine, 10(2):25–??, February 1993. CODEN COMLEF. ISSN 0749-2839.

Plauger:1993:PPc

[Pla93c] P. J. Plauger. Programming on purpose. Computer Language Magazine, 10(3):25–??, March 1993. CODEN COMLEF. ISSN 0749-2839.

Plauger:1993:PPd

[Pla93d] P. J. Plauger. Programming on purpose. Computer Language Magazine, 10(4):25–??, April 1993. CODEN COMLEF. ISSN 0749-2839.

Plauger:1993:PPe

[Pla93e] P. J. Plauger. Programming on purpose. Computer Language Magazine, 10(5):25–??, May 1993. CO-DEN COMLEF. ISSN 0749-2839.

Plauger:1993:PPf

[Pla93f] P. J. Plauger. Programming on purpose. Computer Language Magazine, 10(6):25–??, June 1993. CO-DEN COMLEF. ISSN 0749-2839.

Plum:1991:BSI

[Plu91] Thomas Plum. Building a standard is hard; testing it is just as difficult. Computer Language Magazine, 8(5):38–??, May 1991. CO-DEN COMLEF. ISSN 0749-2839.

Plum:1993:CSP

[Plu93] Tom Plum. Conformance and support for portable code. Computer Language Magazine, 10(5):95–??, May 1993. CODEN COMLEF. ISSN 0749-2839.

Pomerantz:1990:FGC

[Pom90] Dave Pomerantz. A few good colors. Computer Language Magazine, 7(8):32–??, August 1990. CO-DEN COMLEF. ISSN 0749-2839.

Powell:1990:BFV

[Pow90] Jim Powell. B-Tree Filer v. 5.0.

Computer Language Magazine, 7 (1):99–??, January 1990. CODEN COMLEF. ISSN 0749-2839.

Powell:1992:RV

[Pow92a] James E. Powell. Realizer v. 1.0.

Computer Language Magazine, 9 (3):81–??, March 1992. CODEN COMLEF. ISSN 0749-2839.

Powell:1992:SBP

[Pow92b] Jim Powell. Stony Brook Pascal + v. 6.0. Computer Language Magazine, 9(1):85–??, January 1992. CODEN COMLEF. ISSN 0749-2839.

Rabin:1992:RO

[Rab92] Steve Rabin. Reengineering opportunities. Computer Language Magazine, 9(4):51–??, April 1992. CO-DEN COMLEF. ISSN 0749-2839.

Rabin:1993:AUU

[Rab93] Steve Rabin. ANSI 85: Understanding and using the latest COBOL standard. Computer Language Magazine, 10(3):67–??, March 1993. CODEN COMLEF. ISSN 0749-2839.

Rajput:1992:AOP

[Raj92] Wasim E. Rajput. Architecting OS/2 presentation manager applications. Computer Language Magazine, 9(2):61–??, February 1992. CODEN COMLEF. ISSN 0749-2839.

Ravitz:1990:FL

[Rav90] Cary Ravitz. Fancy lines. Computer Language Magazine, 7(8):71– ??, August 1990. CODEN COM-LEF. ISSN 0749-2839.

Reed:1992:ODO

[Ree92] Harvey Reed. Object-oriented design by orthogonality. Computer Language Magazine, 9(1):51– ??, January 1992. CODEN COM-LEF. ISSN 0749-2839.

Reilly:1992:OOC

[Rei92] Douglas J. Reilly. Operator overloading in C++. Computer Language Magazine, 9(10):57–??, October 1992. CODEN COMLEF. ISSN 0749-2839.

Reilly:1993:OD

[Rei93] Douglas Reilly. Object-oriented databases. Computer Language Magazine, 10(1):73–??, January 1993. CODEN COMLEF. ISSN 0749-2839.

Rex:1992:RR

[Rex92] John Rex. Recursion revisited.

Computer Language Magazine, 9 (6):59–??, June 1992. CODEN COMLEF. ISSN 0749-2839.

Reynolds:1992:PCC

[Rey92] Carl W. Reynolds. Polygon clipping: Creating a polygon-fill function. Computer Language Magazine, 9(8):36–??, August 1992. CO-DEN COMLEF. ISSN 0749-2839.

Reynolds:1993:SPC

[Rey93] Carl W. Reynolds. The Sutherland-Hodgman polygon clipping algorithm. Computer Language Magazine, 10(2):51–??, February 1993. CODEN COMLEF. ISSN 0749-2839.

Ribar:1990:OPG

[Rib90] L. John Ribar. OS/2 programming goes Modula-2. Computer Language Magazine, 7(11):83–??, November 1990. CODEN COM-LEF. ISSN 0749-2839.

Riehle:1991:GMO

[Rie91a] Richard Riehle. Generalized methods in object-based languages. Computer Language Magazine, 8 (10):67–??, October 1991. CODEN COMLEF. ISSN 0749-2839.

Riehle:1991:SEC

[Rie91b] Richard Riehle. Software engineering… in COBOL? Computer Language Magazine, 8(1):55–??, January 1991. CODEN COMLEF. ISSN 0749-2839.

Rodini:1990:CWS

[Rod90] Bob Rodini. Crafting WinApps with STDs. Computer Language Magazine, 7(3):45–??, March 1990. CODEN COMLEF. ISSN 0749-2839.

Rogerson:1990:GSW

[Rog90] Dale Rogerson. Getting started with Windows. Computer Language Magazine, 7(3):77–??, March 1990. CODEN COMLEF. ISSN 0749-2839.

Rovira:1990:RD

[Rov90] Charles-A. Rovira. Rovira diagrams. Computer Language Magazine, 7(1):59–??, January 1990. CODEN COMLEF. ISSN 0749-2839.

Saks:1990:CAS

[Sak90a] Dan Saks. C and the ANSI Standard. Computer Language Magazine, 7(5):65–??, May 1990. CO-DEN COMLEF. ISSN 0749-2839.

Saks:1990:MCC

[Sak90b] Dan Saks. MS-DOS for C++ compilers. Computer Language Magazine, 7(10):91–??, October 1990. CODEN COMLEF. ISSN 0749-2839.

Schneier:1992:DIR

[Sch92a] Bruce Schneier. Doing it randomly: Probabilistic algorithms in programming. Computer Language Magazine, 9(8):69–??, August 1992. CODEN COMLEF. ISSN 0749-2839.

Schneier:1992:FMM

[Sch92b] Bruce Schneier. Fast matrix multiplies. Computer Language Magazine, 9(11):49–??, November 1992. CODEN COMLEF. ISSN 0749-2839.

Schneier:1992:SSA

[Sch92c] Bruce Schneier. Sharing secrets among friends. Computer Language Magazine, 9(4):57–??, April 1992. CODEN COMLEF. ISSN 0749-2839.

Sharp:1992:PPT

[Sha92a] Oliver Sharp. Perl: The programmers’s toolbox. Computer Language Magazine, 9(12):40–??, December 1992. CODEN COMLEF. ISSN 0749-2839.

Shaver:1992:CCT

[Sha92b] Dave Shaver. Code comprehension tools. Computer Language Magazine, 9(9):81–??, September 1992. CODEN COMLEF. ISSN 0749-2839.

Skjellum:1984:OPP

[Skj84] Anthony Skjellum. Occam: a parallel processing language from the UK. Computer Language Magazine, 1(3):55–60, November 1984. CODEN COMLEF. ISSN 0749-2839.

Shlaer:1990:RD

[SM90] Sally Shlaer and Stephen J. Mellor. Recursive design. Computer Language Magazine, 7(3):53– ??, March 1990. CODEN COM-LEF. ISSN 0749-2839.

Sperry:1991:FWF

[Spe91] Tyler Sperry. A few words on Forth. Computer Language Magazine, 8(3):55–??, March 1991. CO-DEN COMLEF. ISSN 0749-2839.

Sperry:1992:FF

[Spe92] Tyler Sperry. Four Forths. Computer Language Magazine, 9(1):77– ??, January 1992. CODEN COM-LEF. ISSN 0749-2839.

Steele:1986:CLC

[Ste86] Guy Lewis Steele. Common LISP, and the connection machine. Computer Language Magazine, 11(8): ??, August 1986. CODEN COM-LEF. ISSN 0749-2839.

Stitt:1990:CUI

[Sti90a] Martin Stitt. Capturing user input. Computer Language Magazine, 7(4):61–??, April 1990. CO-DEN COMLEF. ISSN 0749-2839.

Stitt:1990:P

[Sti90b] Martin Stitt. Up periscope! Computer Language Magazine, 7(2): 101–??, February 1990. CODEN COMLEF. ISSN 0749-2839.

Suydam:1990:SPE

[Suy90] Bill Suydam. Smalltalk as a programming environment. Computer Language Magazine, 7(4):34– ??, April 1990. CODEN COMLEF. ISSN 0749-2839.

Sullo:1992:FW

[SW92] Gary Sullo and Mike Williams.

Formal where? Computer Language Magazine, 9(7):51–??, July 1992. CODEN COMLEF. ISSN 0749-2839.

Swanke:1991:RPS

[Swa91a] John E. Swanke. The realities of portable software. Computer Language Magazine, 8(11):57–??, November 1991. CODEN COM-LEF. ISSN 0749-2839.

Swanke:1991:ARE

[Swa91b] Robert S. Swanke. The art of reverse engineering. Computer Language Magazine, 8(6):57–??, June 1991. CODEN COMLEF. ISSN 0749-2839.

Taylor:1986:WAM

[Tay86] Dave Taylor. Wordz that almost match. Computer Language Magazine, 3(11):47–59, November 1986. CODEN COMLEF. ISSN 0749-2839.

Taylor:1990:MMP

[Tay90a] Dave Taylor. Metrowerks Modula-2 PSE. Computer Language Magazine, 7(11):91–??, November 1990. CODEN COMLEF. ISSN 0749-

Taylor:1990:PPE

[Tay90b] Dave Taylor. Professional programmer’s extender. Computer Language Magazine, 7(6):102–??, June 1990. CODEN COMLEF. ISSN 0749-2839.

Taylor:1990:PM

[Tay90c] Dave Taylor. Prolcon for the Mac. Computer Language Magazine, 7 (9):102–??, September 1990. CO-DEN COMLEF. ISSN 0749-2839.

Taylor:1990:UCO

[TH90] David K. Taylor and Alan Hecht. Using CASE for object-oriented design with C++. Computer Language Magazine, 7(11):49–??, November 1990. CODEN COM-LEF. ISSN 0749-2839.

Toutonghi:1990:CA

[Tou90] Michael Toutonghi. 21st century assembler. Computer Language Magazine, 7(6):65–??, June 1990. CODEN COMLEF. ISSN 0749-2839.

Turpin:1992:BES

[Tur92] Bill Turpin. Bringing expert systems to the desktop. Computer Language Magazine, 9(5):71– ??, May 1992. CODEN COMLEF. ISSN 0749-2839.

Tylutki:1990:WGI

[Tyl90] George Tylutki. What good is self? Computer Language Magazine, 7 (11):77–??, November 1990. CO-DEN COMLEF. ISSN 0749-2839.

Veale:1991:PPV

[Vea91a] David Veale. Professional Pascal v. 2.8. Computer Language Magazine, 8(12):97–??, December 1991. CODEN COMLEF. ISSN 0749-2839.

Veale:1991:TPW

[Vea91b] David Veale. Turbo Pascal for Windows. Computer Language Magazine, 8(8):82–??, August 1991. CODEN COMLEF. ISSN 0749-2839.

Veale:1992:CV

[Vea92] David Veale. CommonBase v. 1.2.

Computer Language Magazine, 9 (6):97–??, June 1992. CODEN COMLEF. ISSN 0749-2839.

Wagner:1993:BIO

[Wag93] Richard Wagner. Borland International’s ObjectVision. Computer Language Magazine, 10(4):81–??, April 1993. CODEN COMLEF. ISSN 0749-2839.

Walsh:1993:DSQ

[Wal93] James Walsh. Determining software quality. Computer Language Magazine, 10(4):57–??, April 1993. CODEN COMLEF. ISSN 0749-2839.

Weeks:1990:CGE

[Wee90] Kevin Weeks. Classy graphics:

Encapsulating the Borland Graphics Interface. Computer Language Magazine, 7(10):49–??, October 1990. CODEN COMLEF. ISSN 0749-2839.

Weeks:1992:IYC

[Wee92] Kevin Weeks. Is your code done yet? Computer Language Magazine, 9(4):63–??, April 1992. CO-DEN COMLEF. ISSN 0749-2839.

Weller:1992:BTS

[Wel92] James Weller. Basing tomorrow’s strategies on yesterday’s data. Computer Language Magazine, 9 (3):69–??, March 1992. CODEN COMLEF. ISSN 0749-2839.

Whitney:1990:UTM

[Whi90] Jack Whitney. Unknown text modes. Computer Language Magazine, 7(8):63–??, August 1990. CO-DEN COMLEF. ISSN 0749-2839.

White:1993:HCS

[Whi93] Mitchell White. How to copyright software. Computer Language Magazine, 10(3):82–??, March 1993. CODEN COMLEF. ISSN 0749-2839.

Wiegers:1993:ISE

[Wie93] Karl E. Wiegers. Implementing software engineering in a small software group. Computer Language Magazine, 10(6):55–??, June 1993. CODEN COMLEF. ISSN 0749-2839.

Williams:1990:SSC

[Wil90] Kent Williams. Smart searching with C++. Computer Language Magazine, 7(9):45–??, September 1990. CODEN COMLEF. ISSN 0749-2839.

Winer:1990:BB

[Win90] Ethan Winer. Beyond the basics. Computer Language Magazine, 7 (2):89–??, February 1990. CODEN COMLEF. ISSN 0749-2839.

Wasserman:1991:OSD

[WP91] Anthony I. Wasserman and Peter A. Pircher. Object-oriented structured design and C++. Computer Language Magazine, 8(1):41– ??, January 1991. CODEN COM-LEF. ISSN 0749-2839.

Zeichick:1991:HFI

[Zei91a] Alan Zeichick. How far is Berlin?

Computer Language Magazine, 8 (9):69–??, September 1991. CO-DEN COMLEF. ISSN 0749-2839.

Zeichick:1991:WBS

[Zei91b] Alan Zeichick. Window Builder for Smalltalk/V Windows v. 1.0. Computer Language Magazine, 8 (10):19–??, October 1991. CODEN COMLEF. ISSN 0749-2839.

Zeichick:1992:WGI

[Zei92] Alan L. Zeichick. Watcom Group Inc.’s FORTRAN 77/386 v. 8.5. Computer Language Magazine, 9 (3):17–??, March 1992. CODEN COMLEF. ISSN 0749-2839.